KjSaif:

Graph the ordered pairs for y = 3x + 3 using x={−2,  1, 2}.

KjSaif:

http://prnt.sc/ddt2li

WillieXO:

You can substitute the values of x in the expression to find the values of y (the complementary of the ordered pairs) Let's start with when x =-2 Y = 3(-2) + 3 Y = -3 Thus the 1st ordered pair is (-2,-3) The 2nd when x= 1 Y= 3(1) + 3 Y = 6 (1,6) The 3rd Is when x= 2 Y = 3(2) + 3 Y = 9 (2,9) Plug those coordinates and you'll get your graph

KjSaif:

bank you

KjSaif:

xD

WillieXO:

You welcome XD. One question tho how did you made it to level 100!!!

KjSaif:

i stayed here for months xD

Shadow:

Doing what, I wonder..

WillieXO:

U stayed here!! I thought the site was opened recently mate

WillieXO:

Like 6-7 months ago ult told me.. Besides even 8 months wouldn't get you to level 100 that fast apart from the lack of questions in here and getting medals is difficult..!! That's debatable

563blackghost:

Someone is quite perspective ;)

WillieXO:

Lol ghosty ;)

Ultrilliam:

the site has had bursts of activity... look at the closed questions